# From Lengthy to Lucid: A Systematic Literature Review on NLP Techniques for Taming Long Sentences

### v1

@@ Abstract.
Long sentences have been a persistent issue in written communication for many years since they make it challenging for readers to grasp the main points or follow the initial intention of the writer. This survey, conducted using the PRISMA guidelines, systematically reviews two main strategies for addressing the issue of long sentences: a) sentence compression and b) sentence splitting. An increased trend of interest in this area has been observed since 2005, with significant growth after 2017. Current research is dominated by supervised approaches for both sentence compression and splitting. Yet, there is a considerable gap in weakly and self-supervised techniques, suggesting an opportunity for further research, especially in domains with limited data. In this survey, we categorize and group the most representative methods into a comprehensive taxonomy. We also conduct a comparative evaluation analysis of these methods on common sentence compression and splitting datasets. Finally, we discuss the challenges and limitations of current methods, providing valuable insights for future research directions. This survey is meant to serve as a comprehensive resource for addressing the complexities of long sentences. We aim to enable researchers to make further advancements in the field until long sentences are no longer a barrier to effective communication.
sentence compression, sentence summarization, sentence splitting, split and rephrase, long sentences, systematic review
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXccs: General and reference Surveys and overviewsccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generation
1. Introduction
@@
Several studies going back more than 100 years, have shown that the length of a sentence in words is negatively correlated with its readability (Dubay, 2006). The Oxford guide to plain English recommends sentences of 15 to 20 words (Cutts, 2007). According to the cognitive load theory (Sweller, 2011), humans’ limited working memory capacity makes it difficult to process longer and more complex sentences. Such sentences make it difficult for readers to comprehend the meaning and intention of the writer.
@@
However, people tend to include long and complex sentences in their writing. Therefore, there is an increasing need for automated systems that can effectively deconstruct long sentences into shorter and more comprehensive ones. This is critical in many societal applications such as social media, marketing as well as accessibility services, where long sentences make the content much harder to digest. In addition, shorter sentences can benefit people with reading disabilities and improve many natural language processing (NLP) applications, such as text summarization (Lin, 2003; Zajic et al., 2008; Li et al., 2013), headline (Rush et al., 2015; Filippova et al., 2015) and subtitle generation (Luotolahti and Ginter, 2015), machine translation (Li et al., 2020a) and speech processing (Wu et al., 2007; Wang et al., 2012).
@@
Transforming a long sentence into a more concise form without losing any important information is not always a straightforward task. There are two ways for addressing this problem: a) shortening the long sentence, without losing key information from it, and b) splitting it into two or more short sentences. The first task, which is met in the literature as sentence compression or sentence summarization receives a long sentence as input and generates a shortened version retaining the most important information. Sentence compression can be extractive or abstractive. Extractive methods work by simply deleting words from the input sentence in contrast to abstractive ones which involve more complex paraphrasing, rewriting, and reordering operations. Studies of human summarizers have shown that during compression, besides deleting words, they commonly also do paraphrasing, generalization, and reordering (Jing, 2000). On the other hand, human-written reference summaries have been found to exhibit high word overlap with the source sentence, even preserving word order to a large extent (Schumann et al., 2020). Sentence splitting segments a given long sentence into two or more shorter sentences of equivalent meaning. We can also meet sentence splitting in the literature as split and rephrase task (Narayan et al., 2017; Aharoni and Goldberg, 2018; Zhang et al., 2020b; Kim et al., 2021). The split and rephrase task does not involve any deletion or simplification, apart from the required syntax modification of the shorter sentences. In many prior studies, sentence splitting is typically met as a preprocessing step within machine translation (Lee et al., 2008), discourse relation (Yu et al., 2019) or sentence simplification systems (Siddharthan, 2014).
@@
Addressing long sentences has been a subject of research in the literature for over 20 years, as shown in Fig. 1, where the number of publications related to sentence compression and splitting from 2000 to 2022 are presented, as discovered through our systematic survey methodology (see Section 2). We can see that there has been a growing interest in this subject since 2005, which has intensified after 2017. Despite the increasing research interest, there are limited surveys in the field. Those are mainly focused on extractive sentence compression (Fathima and Raseek, 2018) and evaluation of sentence compression methods (Napoles et al., 2011b). This paper provides a comprehensive review that explores both sentence compression and sentence splitting for resolving long sentences. To the best of our knowledge, this is the first systematic review on this topic.
Figure 1. Moving average of publications per year from 2000 to 2022 for sentence compression and sentence splitting.
@@
  • It provides a systematic review of state-of-the-art methods for dealing with long sentences, including both sentence compression and sentence splitting.
  • It categorizes the main approaches with the most representative methods into a comprehensive taxonomy.
  • It provides a comparative analysis between existing methods on widely used datasets for this task.
  • It discusses the challenges and limitations of existing methods.
@@
2. Survey Methodology
This survey was conducted following the PRISMA guidelines (Page et al., 2021) for reporting systematic reviews. In this section, we provide an overview of the survey methodology, including eligibility criteria, information sources, and search strategy, as well as the PRISMA flowgram and the results of the study selection.
All original works concerning the task of dealing with long sentences were considered eligible for this systematic review. The inclusion criteria adopted were: i) publication date between 2000 and 2023; ii) written in the English language; iii) works providing a full-text article; and iv) journal and conference publications. Unpublished works such as dissertation studies, theses and technical reports were excluded from this review. Due to the large number of works, studies with few citations (less than 10 citations in total) that were published prior to 2017 were also excluded from this review.
@@ 2. Survey Methodology
The following search query was used to identify relevant papers: (“sentence summarization” OR “sentence compression” OR “sentence splitting” OR “split and rephrase”). We arrived at this query by starting from (‘sentence compression” OR “sentence splitting”) and gradually expanding it with additional synonymous keywords that we discovered on the way. We limit our search to the first 300 records of each database to maintain a manageable result pool. After the initial search, a total of 1,479 records were collected via digital databases (Google Scholar n = 300, Scopus n = 300, ACM Digital Library n = 300, Springer Digital Library n = 300, IEEE Explore Digital Library n = 121, ArXiv n = 94) and hand search reference lists (n = 64). 305 records were removed due to duplication before the screening stage. Further 955 records were excluded during the title and abstract screening since they did not fall within the scope of this review. After reviewing the full-text, records were excluded for eligibility due to language issues (n = 3), publication date (n = 1), few citations (n = 95) and unpublished works (n = 10). During the study selection, 110 records fully met the initial criteria and are included in this review. The study selection flowgram is shown in detail in Fig. 2.
Figure 2. PRISMA flow diagram of the study selection process.
3. Sentence Compression
@@
In this section, we review methods, datasets, and evaluation metrics for the task of sentence compression. In addition, we provide a comparative evaluation of the most representative methods for this task. Finally, we discuss other trends observed in the literature including length control and the increasingly relevant domain of multilingual and cross-lingual sentence compression.
Figure 3. Taxonomy of sentence compression methods
3.1. Methods
@@
Sentence compression can be either extractive or abstractive. Extractive methods reduce the length of the input sentence by simply removing words while abstractive methods achieve compression using additional operations besides deleting such as paraphrasing, rewriting and reordering. However, both approaches share the same underlying techniques (e.g., statistical methods, reinforcement learning, deep learning). Thus, this review categorizes the methods based on the employed technique. Figure 4 shows the number of publications from 2000 to 2022, along with landmarks of key technology adoption. Early works from 2000 to 2005 were limited and primarily relied on statistical methods. However, a significant increase in interest became evident after 2008, especially from 2015 onwards with the introduction of deep learning techniques. A spike in interest is also observed after the emergence of Transformers. The increasing attention in sentence compression methods can be attributed to the remarkable advancements in the field of NLP, as well as the ever increasing textual information overload from various sources. A high-level taxonomy of sentence compression methods is shown in Figure 3.
@@
Figure 4. Moving average of publications per year from 2000 to 2022 for sentence compression along with landmarks of key technology adoption.
3.1.1. Statistical Methods
@@
Early works on extractive sentence compression involve the use of statistical methods like noisy channels and maximum entropy models. Such methods aim to estimate the joint probability distribution $P(x,y)$ of a given long sentence $x$ and its target compression $y$. Noisy channel models are used to determine which words must be removed from the input sentence (Knight and Marcu, 2000, 2002a; Turner and Charniak, 2005; Galley and McKeown, 2007). These models work by adding words to the original short sentence resulting to a noisy long sentence. The goal is to find the compression that maximizes the probability of the short sentence to be the condensed version of the long sentence. Knight and Marcu (2000) are the first to employ a noisy channel with a small dataset of 1,067 compressions. Turner and Charniak (2005) adapt this model for unsupervised learning without the need of parallel data, while other approaches involve additional lexical variables into the model (Galley and McKeown, 2007) or extend it with a maximum entropy model (Unno et al., 2006).
3.1.2. Optimization Techniques
@@
Several techniques approach the task of sentence compression as an optimization problem to find the optimal solution or set of solutions from a given set of possible solutions, subject to defined constraints. These approaches include Integer Linear Programming (ILP) (Clarke and Lapata, 2006a, 2007, 2008, 2010; Thadani and McKeown, 2013) , Dynamic Programming (DP) (Napoles et al., 2011a; Liu et al., 2022b) or other optimization algorithms such as first-choice hill climbing (Schumann et al., 2020; Liu et al., 2022a).
@@
Early methods formulate the task of sentence compression as a constrained optimization problem with ILP. ILP is an optimization technique where a linear objective function and decision variables are subject to linear constraints that are represented as linear inequalities. Similar to other optimization problems, ILP aims to find a global optimum solution by searching a given space of possible solutions. The decision variables must be integer values while the objective function is typically formulated as a weighted sum of the decision variables. For example, Clarke and Lapata (2006a) keep the words of the input sentence that maximize a language model’s objective function. The scoring function is subject to several hand-crafted grammar and syntactic constraints. Clarke and Lapata (2007) exploit the discourse relations between sentences in the whole input document instead of handling each sentence separately. Clarke and Lapata (2008) set global constraints to find the optimal compression. Unlike the local constraints of the previous approaches, global constraints allow for considering the whole sentence and its relations rather than processing isolated segments and units. This approach includes both supervised, semi-supervised and unsupervised models. Thadani and McKeown (2013) use commodity flow constraints that allow for changing the word ordering to ensure the appropriate syntax of the compression, unlike the hard global constraints of (Clarke and Lapata, 2008) that are restricted to the same token ordering as the input sentence. Thadani (2014) improves the latency of previous methods using an approximate inference algorithm with Lagrangian relaxation.
@@
Other approaches employ DP to break down the task of sentence compression into smaller sub-problems. The final solution is being constructed by combining the solutions to these sub-problems. More specifically, Napoles et al. (2011a) introduce a method where paraphrasing and dynamic programming are combined for character-based compression. This method works by replacing existing words in the input with their paraphrases to extract the optimal solution of non-overlapping paraphrases that minimizes the character length of the sentence. Similarly, in (Liu et al., 2022b) the length reduction is imposed as a character-level length constraint, with possible lengths to be divided into buckets.
@@
Recent methods employ a first-choice hill climbing algorithm that starts with an initial candidate solution and iteratively explores local solutions to find a better one. More specifically, Schumann et al. (2020) introduce an objective function that is based on hill climbing for removing words from the initial sentence. This function is used to generate extractive compressions in an unsupervised way. In (Liu et al., 2022a) the same idea is adopted for generating pseudo-compressions, while a Transformer model is trained on these compressions.
3.1.3. Dependency Parsing
@@
Dependency parsing is the task of extracting syntactic and grammatical relations between each word in a sentence to better analyze its structure. Dependency parsing is widely used to solve the task of sentence compression in both extractive (Filippova and Strube, 2008; Filippova and Altun, 2013; Galanis and Androutsopoulos, 2010) and abstractive settings (Cohn and Lapata, 2008, 2009; Galanis and Androutsopoulos, 2011).
@@
Extractive approaches employ a dependency tree to prune unimportant words of the input sentence (Filippova and Strube, 2008; Filippova and Altun, 2013; Galanis and Androutsopoulos, 2010). For example, Filippova and Strube (2008) introduce an unsupervised method for sentence compression based on a dependency tree by pruning sub-trees of the original sentence that do not contain important details. Filippova and Altun (2013) propose a supervised method based on dependency parsing that builds upon previous unsupervised work (Filippova and Strube, 2008). Instead of using pre-computed edge weights, this method employs a linear function over different feature representations (syntactic, structural, semantic, and lexical) to enhance the readability and informativeness of the compression. Galanis and Androutsopoulos (2010) employ a dependency tree to generate different compression candidates by pruning unimportant words from the source sentence, while a support vector regression (SVR) model is used to select the best candidate. This method can successfully perform extractive compression without the need for defined rules, as in (Clarke and Lapata, 2008).
@@
Cohn and Lapata (2008) propose the first abstractive sentence compression model that employs a tree-to-tree rewriting task with additional rewriting operations such as substitution, insertion and reordering, instead of simple deletions. Cohn and Lapata (2009) approach this task using a syntactic dependency tree with a large set of features.  Galanis and Androutsopoulos (2011) introduce an abstractive approach with a generate-and-rank sentence model. First, different extractive candidates are generated using a dependency tree, which are then transformed into abstractive candidates with paraphrasing tools. The best candidate is selected based on an SVR ranking module. Similarly, Cohn and Lapata (2013) build upon rewriting and paraphrasing operations for abstractive sentence compression rather than simply deleting words from the input.
@@
Even though these approaches can be successful in generating grammatically correct sentences, they might result in pruning important words from the input sentence. Recent approaches (Kamigaito et al., 2018; Kamigaito and Okumura, 2020) claim that a long distance between the root node and the child nodes in a dependency tree should not be neglected. For example, Kamigaito et al. (2018) use an LSTM paired with a dependency parser to keep words that belong to the parent dependency chain of the word that should be retained.  Kamigaito and Okumura (2020) extend this approach to consider the dependency chain of the children as well. Another approach that employs dependency parsing in combination with neural architectures is introduced in (Park et al., 2021). More specifically, in this work, a neural model for extractive sentence compression is proposed, where the dependency tree of each word of the input sentence is encoded using a graph convolutional network (GCN). Then, a pre-trained BERT model is combined with multiple GCN layers, while a scoring layer decides which words of the input sentence should be retained or removed.
3.1.4. Deep Learning

### v2

@@ Abstract
Long sentences have been a persistent issue in written communication for many years since they make it challenging for readers to grasp the main points or follow the initial intention of the writer. This survey, conducted using the PRISMA guidelines, systematically reviews two main strategies for addressing the issue of long sentences: a) sentence compression and b) sentence splitting. An increased trend of interest in this area has been observed since 2005, with significant growth after 2017. Current research is dominated by supervised approaches for both sentence compression and splitting. Yet, there is a considerable gap in weakly and self-supervised techniques, suggesting an opportunity for further research, especially in domains with limited data. We also observe that despite their potential, Large Language Models (LLMs) have not yet been widely explored in this area. In this survey, we categorize and group the most representative methods into a comprehensive taxonomy. We also conduct a comparative evaluation analysis of these methods on common sentence compression and splitting datasets. Finally, we discuss the challenges and limitations of current methods, providing valuable insights for future research directions. This survey is meant to serve as a comprehensive resource for addressing the complexities of long sentences. We aim to enable researchers to make further advancements in the field until long sentences are no longer a barrier to effective communication.
1 Introduction
@@
Several studies going back more than 100 years, have shown that the length of a sentence in words is negatively correlated with its readability Dubay2006TheStudies . The Oxford guide to plain English recommends sentences of 15 to 20 words Cutts2007OxfordEnglish . According to the cognitive load theory sweller2011cognitive , humans’ limited working memory capacity makes it difficult to process longer and more complex sentences. Such sentences make it difficult for readers to comprehend the meaning and intention of the writer.
@@
However, people tend to include long and complex sentences in their writing. Therefore, there is an increasing need for automated systems that can effectively deconstruct long sentences into shorter and more comprehensive ones. This is critical in many societal applications such as social media, marketing as well as accessibility services, where long sentences make the content much harder to digest. In addition, shorter sentences can benefit people with reading disabilities and improve many natural language processing (NLP) applications, such as text summarization lin2003improving ; zajic2008 ; li2013document , headline Rush2015ASummarization ; Filippova2015 and subtitle generation luotolahti2015sentence , machine translation li2020explicit and speech processing wu2007 ; wang2012two .
@@
Transforming a long sentence into a more concise form without losing any important information is not always a straightforward task. There are two ways for addressing this problem: a) shortening the long sentence, without losing key information from it, and b) splitting it into two or more short sentences. The first task, which is met in the literature as sentence compression or sentence summarization receives a long sentence as input and generates a shortened version retaining the most important information. Sentence compression can be extractive or abstractive. Extractive methods work by simply deleting words from the input sentence in contrast to abstractive ones which involve more complex paraphrasing, rewriting, and reordering operations. Studies of human summarizers have shown that during compression, besides deleting words, they commonly also do paraphrasing, generalization, and reordering Jing2000 . On the other hand, human-written reference summaries have been found to exhibit high word overlap with the source sentence, even preserving word order to a large extent Schumann2020DiscreteExtraction . Sentence splitting segments a given long sentence into two or more shorter sentences of equivalent meaning. We can also meet sentence splitting in the literature as split and rephrase task narayan-etal-2017-split ; aharoni-goldberg-2018-split ; Zhang2020 ; kim-etal-2021-bisect . The split and rephrase task does not involve any deletion or simplification, apart from the required syntax modification of the shorter sentences. In many prior studies, sentence splitting is typically met as a preprocessing step within machine translation lee2008transformation , discourse relation yu2019gumdrop or sentence simplification systems siddharthan2014survey .
@@
Addressing long sentences has been a subject of research in the literature for over 20 years, as shown in Fig. 1, where the number of publications related to sentence compression and splitting from 2000 to 2025 are presented, as discovered through our systematic survey methodology (see Section 2). We can see that there has been a growing interest in this subject since 2005, which has intensified after 2017. Despite the increasing research interest, there are limited surveys in the field. Those are mainly focused on extractive sentence compression fathima2018review and evaluation of sentence compression methods napoles2011evaluating . This paper provides a comprehensive review that explores both sentence compression and sentence splitting for resolving long sentences. To the best of our knowledge, this is the first systematic review on this topic.
Figure 1: Cumulative number of publications from 2000 to 2025 for sentence compression and sentence splitting, demonstrating the momentum and growing interest in the field.
@@
  • We provide a systematic review of state-of-the-art methods for dealing with long sentences, including both sentence compression and sentence splitting.
  • We categorize the main approaches with the most representative methods into a comprehensive taxonomy.
  • We provide a comparative analysis between existing methods on widely used datasets for this task.
  • We discuss the challenges and limitations of existing methods.
@@
2 Survey Methodology
This survey was conducted following the PRISMA guidelines page2021prisma for reporting systematic reviews. In this section, we provide an overview of the survey methodology, including eligibility criteria, information sources, and search strategy, as well as the PRISMA flowgram and the results of the study selection.
All original works concerning the task of dealing with long sentences were considered eligible for this systematic review. The inclusion criteria adopted were: i) publication date between 2000 and 2025; ii) written in the English language; iii) works providing a full-text article; and iv) journal and conference publications. Unpublished works such as dissertation studies, theses and technical reports were excluded from this review. Due to the large number of works, studies with few citations (less than 10 citations in total) that were published prior to 2017 were also excluded from this review.
@@ 2 Survey Methodology
The following search query was used to identify relevant papers: (“sentence summarization” OR “sentence compression” OR “sentence splitting” OR “split and rephrase”). We arrived at this query by starting from (‘sentence compression” OR “sentence splitting”) and gradually expanding it with additional synonymous keywords that we discovered on the way. We initially searched using all the specified criteria for the years 2000 to 2022, limiting our search to the first 300 records in each database to maintain a manageable result pool. We then repeated this process for the years 2023 to 2025, limiting our search to 30 records per database, as no relevant results were found beyond this threshold.
@@
A total of 1,633 records were collected via digital databases (Google Scholar n = 330, Scopus n = 330, ACM Digital Library n = 330, Springer Digital Library n = 330, IEEE Explore Digital Library n = 129, ArXiv n = 120) and hand search reference lists (n = 64). 349 records were removed due to duplication before the screening stage. Further 1,043 records were excluded during the title and abstract screening since they did not fall within the scope of this review. After reviewing the full-text, records were excluded for eligibility due to language issues (n = 3), publication date (n = 1), few citations (n = 95) and unpublished works (n = 12). During the study selection, 130 records fully met the initial criteria and are included in this review. The study selection flowgram is shown in detail in Fig. 2.
Figure 2: PRISMA flow diagram of the study selection process.
3 Sentence Compression
@@
In this section, we review methods, datasets, and evaluation metrics for the task of sentence compression. In addition, we provide a comparative evaluation of the most representative methods for this task. Finally, we discuss other trends observed in the literature including length control, the increasingly relevant domain of multilingual and cross-lingual sentence compression and the growing interest in multimodal sentence summarization.
Figure 3: Taxonomy of sentence compression methods
3.1 Methods
@@
Sentence compression can be either extractive or abstractive. Extractive methods reduce the length of the input sentence by simply removing words while abstractive methods achieve compression using additional operations besides deleting such as paraphrasing, rewriting and reordering. However, both approaches share the same underlying techniques (e.g., statistical methods, reinforcement learning, deep learning). Thus, this review categorizes the methods based on the employed technique. Figure 4 shows the cumulative number of publications from 2000 to 2024, along with landmarks of key technology adoption. Early works from 2000 to 2005 were limited and primarily relied on statistical methods. However, a significant increase in interest became evident after 2008, especially from 2015 onwards with the introduction of deep learning techniques. A spike in interest is also observed after the emergence of Transformers. The increasing attention in sentence compression methods can be attributed to the remarkable advancements in the field of NLP, as well as the ever increasing textual information overload from various sources. A high-level taxonomy of sentence compression methods is shown in Figure 3.
@@
Figure 4: Cumulative number of publications per year from 2000 to 2025 for sentence compression along with landmarks of key technology adoption.
3.1.1 Statistical Methods
@@
Early works on extractive sentence compression involve the use of statistical methods like noisy channels and maximum entropy models. Such methods aim to estimate the joint probability distribution $P(x,y)$ of a given long sentence $x$ and its target compression $y$. Noisy channel models are used to determine which words must be removed from the input sentence Knight2000 ; Knight2002 ; turner2005supervised ; Galley2007 . These models work by adding words to the original short sentence resulting to a noisy long sentence. The goal is to find the compression that maximizes the probability of the short sentence to be the condensed version of the long sentence. Knight2000 are the first to employ a noisy channel with a small dataset of 1,067 compressions. turner2005supervised adapt this model for unsupervised learning without the need of parallel data, while other approaches involve additional lexical variables into the model Galley2007 or extend it with a maximum entropy model unno2006trimming .
3.1.2 Optimization Techniques
@@
Several techniques approach the task of sentence compression as an optimization problem to find the optimal solution or set of solutions from a given set of possible solutions, subject to defined constraints. These approaches include Integer Linear Programming (ILP) clarkelapata2006constraint ; clarke2007modelling ; Clarke2008GlobalApproach ; clarkelapata2010discourse ; thadani2013-sentence , Dynamic Programming (DP) napoles2011paraphrastic ; Liu2022Controlling or other optimization algorithms such as first-choice hill climbing Schumann2020DiscreteExtraction ; liu2022learning .
@@
Early methods formulate the task of sentence compression as a constrained optimization problem with ILP. ILP is an optimization technique where a linear objective function and decision variables are subject to linear constraints that are represented as linear inequalities. Similar to other optimization problems, ILP aims to find a global optimum solution by searching a given space of possible solutions. The decision variables must be integer values while the objective function is typically formulated as a weighted sum of the decision variables. For example, clarkelapata2006constraint keep the words of the input sentence that maximize a language model’s objective function. The scoring function is subject to several hand-crafted grammar and syntactic constraints. clarke2007modelling exploit the discourse relations between sentences in the whole input document instead of handling each sentence separately. Clarke2008GlobalApproach set global constraints to find the optimal compression. Unlike the local constraints of the previous approaches, global constraints allow for considering the whole sentence and its relations rather than processing isolated segments and units. This approach includes both supervised, semi-supervised and unsupervised models. thadani2013-sentence use commodity flow constraints that allow for changing the word ordering to ensure the appropriate syntax of the compression, unlike the hard global constraints of Clarke2008GlobalApproach that are restricted to the same token ordering as the input sentence. thadani2014approximation improves the latency of previous methods using an approximate inference algorithm with Lagrangian relaxation.
@@
Other approaches employ DP to break down the task of sentence compression into smaller sub-problems. The final solution is being constructed by combining the solutions to these sub-problems. More specifically, napoles2011paraphrastic introduce a method where paraphrasing and dynamic programming are combined for character-based compression. This method works by replacing existing words in the input with their paraphrases to extract the optimal solution of non-overlapping paraphrases that minimizes the character length of the sentence. Similarly, in Liu2022Controlling the length reduction is imposed as a character-level length constraint, with possible lengths to be divided into buckets.
@@
Recent methods employ a first-choice hill climbing algorithm that starts with an initial candidate solution and iteratively explores local solutions to find a better one. More specifically, Schumann2020DiscreteExtraction introduce an objective function that is based on hill climbing for removing words from the initial sentence. This function is used to generate extractive compressions in an unsupervised way. In liu2022learning the same idea is adopted for generating pseudo-compressions, while a Transformer model is trained on these compressions.
3.1.3 Dependency Parsing
@@
Dependency parsing is the task of extracting syntactic and grammatical relations between each word in a sentence to better analyze its structure. Dependency parsing is widely used to solve the task of sentence compression in both extractive Filippova2008 ; Filippova2013OvercomingCompression ; Galanis2010 and abstractive settings Cohn2008 ; Cohn2009 ; Galanis2011ACompressor .
@@
Extractive approaches employ a dependency tree to prune unimportant words of the input sentence Filippova2008 ; Filippova2013OvercomingCompression ; Galanis2010 . For example, Filippova2008 introduce an unsupervised method for sentence compression based on a dependency tree by pruning sub-trees of the original sentence that do not contain important details. Filippova2013OvercomingCompression propose a supervised method based on dependency parsing that builds upon previous unsupervised work Filippova2008 . Instead of using pre-computed edge weights, this method employs a linear function over different feature representations (syntactic, structural, semantic, and lexical) to enhance the readability and informativeness of the compression. Galanis2010 employ a dependency tree to generate different compression candidates by pruning unimportant words from the source sentence, while a support vector regression (SVR) model is used to select the best candidate. This method can successfully perform extractive compression without the need for defined rules, as in Clarke2008GlobalApproach .
@@
Cohn2008 propose the first abstractive sentence compression model that employs a tree-to-tree rewriting task with additional rewriting operations such as substitution, insertion and reordering, instead of simple deletions. Cohn2009 approach this task using a syntactic dependency tree with a large set of features.  Galanis2011ACompressor introduce an abstractive approach with a generate-and-rank sentence model. First, different extractive candidates are generated using a dependency tree, which are then transformed into abstractive candidates with paraphrasing tools. The best candidate is selected based on an SVR ranking module. Similarly, cohn2013 build upon rewriting and paraphrasing operations for abstractive sentence compression rather than simply deleting words from the input.
@@
Even though these approaches can be successful in generating grammatically correct sentences, they might result in pruning important words from the input sentence. Recent approaches kamigaito2018higher ; kamigaito2020syntactically claim that a long distance between the root node and the child nodes in a dependency tree should not be neglected. For example, kamigaito2018higher use an LSTM paired with a dependency parser to keep words that belong to the parent dependency chain of the word that should be retained.  kamigaito2020syntactically extend this approach to consider the dependency chain of the children as well. Another approach that employs dependency parsing in combination with neural architectures is introduced in park2021sentence . More specifically, in this work, a neural model for extractive sentence compression is proposed, where the dependency tree of each word of the input sentence is encoded using a graph convolutional network (GCN). Then, a pre-trained BERT model is combined with multiple GCN layers, while a scoring layer decides which words of the input sentence should be retained or removed.
3.1.4 Deep Learning